# FinTruthQA: A Benchmark Dataset for Evaluating the Quality of Financial Information Disclosure

### v1

@@ Abstract
Accurate and transparent financial information disclosure is crucial in the fields of accounting and finance, ensuring market efficiency and investor confidence. Among many information disclosure platforms, the Chinese stock exchanges’ investor interactive platform provides a novel and interactive way for listed firms to disclose information of interest to investors through an online question-and-answer (Q&A) format. However, it is common for listed firms to respond to questions with limited or no substantive information, and automatically evaluating the quality of financial information disclosure on large amounts of Q&A pairs is challenging. This paper builds a benchmark FinTruthQA, that can evaluate advanced natural language processing (NLP) techniques for the automatic quality assessment of information disclosure in financial Q&A data. FinTruthQA comprises 6,000 real-world financial Q&A entries and each Q&A was manually annotated based on four conceptual dimensions of accounting: question identification, question relevance, answer readability, and answer relevance. We benchmarked various NLP techniques on FinTruthQA, including statistical machine learning models, pre-trained language model and their fine-tuned versions, as well as the large language model (LLM) GPT-4. Experiments showed that existing NLP models have strong predictive ability for question identification and question relevance tasks, but are suboptimal for answer readability and answer relevance tasks. By establishing this benchmark, we provide a robust foundation for the automatic evaluation of information disclosure, significantly enhancing the transparency and quality of financial reporting. FinTruthQA can be used by auditors, regulators, and financial analysts for real-time monitoring and data-driven decision-making, as well as by researchers for advanced studies in accounting and finance, ultimately fostering greater trust and efficiency in the financial markets.
@@ 1 Introduction
Within the financial market, enhancing transparency and timeliness in information disclosure can mitigate information friction, empower investors to make informed decisions, and foster market efficiency. According to Lee and Zhong (2022), retail investors struggle to integrate the existing public information, and increased online platform activities are associated with more active market transactions, such as greater trading volume, return volatility, and market liquidity. The investor interactive platforms established by the Shanghai Stock Exchange (SSE) [^1] http://www.sse.com.cn/ and Shenzhen Stock Exchange (SZSE) [^2] https://www.szse.cn/index/index.html offer a unique avenue for retail investors to engage directly with listed firms. In contrast to conventional modes of disclosure, such as mandatory disclosures and traditional voluntary disclosures like conference calls, which follow specific procedures and formats, this format enables direct, interactive, near real-time communication between firms and ordinary investors through the format of Q&A. By analyzing the interactions on these platforms, researchers can gain insights into how effectively companies communicate with investors, the types of information that are most valued by investors, and how the disclosure practices of listed firms may impact their corresponding trading patterns in financial markets. This will enhance financial market regulation and provide deeper insights into corporate disclosure practices.
@@
The interactive nature of these platforms provides a rich dataset for assessing the quality of information disclosure. However, the sheer volume of data and the diversity in the quality of questions and responses pose significant challenges for large-scale analysis. Leuz and Wysocki (2016) summarize that the commonly used measures of disclosure in the financial sector primarily capture the quantity rather than the quality of the information provided. However, it is possible to construct the quality dimension of disclosure by focusing on specific disclosures, and Natural Language Processing (NLP) emerges as a promising solution to evaluate the quality of information disclosure at scale. Existing applications of NLP in automatic financial and accounting text analysis have demonstrated success in various domains, such as automatic analysis of annual financial reports(Smailović et al., 2017) and automatic sentiment analysis system on financial news articles(Van de Kauter et al., 2015). However, to the best of our knowledge, there is no standard dataset that evaluates the quality of financial information disclosure in these Q&A platforms.
@@
In this work, we present a Q&A information disclosure quality assessment benchmark that evaluates Chinese Q&A data sourced from interactive platforms for communication between investors and listed companies, as established by SSE and SZSE. In developing this model, we have explored several technical challenges, including the interpretation of natural language, context sensitivity, and the subjective nature of relevance and readability assessments. This paper details the development, implementation, and evaluation of our NLP model, emphasizing its potential impact on the financial sector and the broader implications for automated quality assessment in investor communications. Datasets and code are available at https://github.com/bethxx99/FinTruthQA.
@@ 2 Related work
The measurement issue pertaining to disclosures has been a persistent challenge in empirical accounting research(Leuz and Wysocki, 2016). Predominantly, comprehensive measures have focused on quantifying disclosures, encompassing criteria such as the presence or absence of quarterly and annual reports, as well as numerical counts of activities like conference calls and earnings forecasts(Lee and Zhong, 2022). Another commonly used measure is the Association for Investment Management and Research (AIMR) rankings, based on annual surveys of financial analysts who evaluate U.S. firms’ disclosure practices(Nagar et al., 2003; Healy et al., 1999; Welker, 1995). However, they are limited to large U.S. firms and cover only a specific time period. Furthermore, alternative measures designed for specific disclosures typically employ a keyword-based approach to analyze the tone and sentiment of well-established, traditional disclosures such as earnings forecasts and conference call transcripts(Bochkay et al., 2023; Frankel et al., 2010; Li, 2010).
@@
Financial documents, such as annual reports, financial news articles, and economic policy statements, are typically composed in technical language in a non-structured text format and possess unique features distinct from other domains.(Chang et al., 2016) Recent studies have focused on developing pre-trained models tailored for the financial domain (Araci, 2019; Yang et al., 2020; Liu et al., 2021), with a primary emphasis on the English language. These efforts are predominantly oriented towards sentiment classification tasks in the financial context.(Wan et al., 2021) Additional research in this field has concentrated on question-answering tasks, involving datasets that combine textual and tabular information, where numerical analysis is often essential to derive answers (Zhu et al., 2021; Zhao et al., 2022; Li et al., 2022). To our knowledge, there has been no prior research or dataset dedicated to the automatic quality assessment of financial Q&A data.
@@ 3.1 Data collection
The data for this study was obtained from the interactive platforms for communication between investors and listed companies established by the Shanghai Stock Exchange and the Shenzhen Stock Exchange. We collected over 900,000 Q&A entries, among which we randomly selected 6,000 samples for manual annotation, with 3,000 entries from each exchange’s platform. The data collection process involved using web scraping techniques to extract the Q&A entries from the official platforms. The data was then manually reviewed and processed to ensure quality and consistency. The ethics statement and relevant discussion can be found in Supplement Section A.
Figure 2: Example of Q&A pair and quality evaluation criteria.
@@ 3.2 Data annotation
We focuses on four key information disclosure quality evaluation criteria: question identification, question relevance, answer relevance, and answer readability. These criteria are widely recognized as crucial indicators of information quality and are important for investors to consider when evaluating Q&A information. Specifically, the criteria in FinTruthQA are defined as follows:
@@
    Question identification: This criterion assesses whether the question from an investor is a genuine question, as opposed to a suggestion or complaint. Each question is assigned a binary label: Positive, indicating the entry is a real question, or Negative, indicating the entry is not a real question. Data entries with questions that do not meet the definition of a real question are excluded from further evaluation based on other criteria.
@@
    Question relevance: This criterion evaluates whether a question is pertinent to the investment topic, labeling questions as Positive if they are relevant and Negative if they are not. It assesses whether the question is related to the company’s shareholders, financial indicators, industry conditions, or other relevant investment topics.
@@
    Answer readability: This criterion evaluates whether an answer is easy to understand and well-written. The readability score ranges from 1 to 4, with readability decreasing from 4 to 1. A score of 4 indicates that the answer is very readable, clear, and concise, free of ambiguity. A score of 3 indicates moderate readability with some issues. A score of 2 indicates significant readability issues. A score of 1 indicates the answer is unreadable, with major ambiguities, incoherent sentences, and content unrelated to the question.
@@
    Answer relevance: This criterion evaluates whether an answer provides a relevant and informative response to the question asked. It assesses whether the answer provides enough detail and explanation to address the question. Answers are classified into three categories: 1, indicating the answer is completely unrelated to the question (none of the questions were answered); 2, indicating the answer is partially related (some questions were answered) or difficult to judge; and 3, indicating the answer is fully related to the question (all questions were answered).
The importance of these criteria stems from their collective ability to enhance the quality and reliability of information on investor communication platforms. Annotators with finance and accounting backgrounds were engaged to annotate the Q&A data. Figure 2 illustrates an example of the annotation. A detailed annotation process is listed in the Supplement Section B.
@@ 3.3 Quality control
For the answer relevance task, we observed that the raw annotations included noise due to the subjectivity of this task. Hence, we conducted an additional quality control step to correct the raw annotations. We implemented an enhanced correction process inspired by Confident Learning (Northcutt et al., 2021) principles. Sample data points demonstrating inaccuracies and implementation details are presented in Supplement Section C.
3.4 Data pre-processing
We found that the Shanghai Stock Exchange (SSE) data includes the company name and stock code at the beginning of the question texts, while the Shenzhen Stock Exchange (SZSE) data does not. We compared experiments on Q&A pairs with and without the company name as the prefix, and observed negligible differences. We removed these prefixes from the SSE data in our experiments.
@@ 3.5 Dataset statistics
The complete annotated dataset comprises 6,000 samples labeled for question identification, question relevance, answer relevance, and answer readability tasks, with each data point having four distinct labels corresponding to these tasks. The dataset was randomly divided into training, validation, and testing sets in a ratio of 4:1:1. The label distributions for each task are detailed in Tables 1. The provided statistics reflect the data after undergoing our quality control process. Note that our later analysis focuses exclusively on real questions. Only for the first task (question identification), did we utilize the complete set of 6,000 data samples. In the subsequent three tasks, we removed samples that were not questions. This decision was based on the rationale that non-question data could skew the model’s understanding and performance, especially for tasks like question relevance, where the context of a real question is crucial for accurate assessment. It is worth noticing that the labels of task question relevance are highly imbalanced, with only 6 and 7 negative labels in the validation and test dataset, respectively. We decided to keep this task as this reflects the real-world data distribution, and the concept of question relevance is important in financial information disclosure. Figure  3 shows the length distribution of questions and answers in characters of all the 6,000 QA pairs.
@@
Task Label Train Validation Test Label_Num Freq Total
Question identification Negative 327 90 81 498 8.30% 6000
Positive 3673 910 919 5502 91.70%
Question relevance Negative 714 6 7 27 0.49% 5502
Positive 3654 911 910 5475 99.51%
Answer readability 1 106 190 198 1094 19.88% 5502
@@ 4.1 Model selection
To fully evaluate the performance of different methods in FinTruthQA, we benchmark several models that cover Machine Learning (ML)-based, Pre-trained Language Model (PLM)-based, PLM with continued Pre-Taining (CPT)-based, and LLM-based. Given the distinct nature and varying difficulty of different tasks, we applied different approaches accordingly.
@@
ML-based. We modeled these tasks as text classification and employed various advanced machine learning models, including Logistic Regression (LR)(Hosmer Jr et al., 2013), Support Vector Machine (SVM)(Hearst et al., 1998), K-nearest neighbor (KNN)(Peterson, 2009), and Random Forest (RF)(Breiman, 2001). For semantic feature extraction, we first utilized the Jieba toolkit  [^3] https://github.com/fxsjy/jieba for Chinese word segmentation and then extracted the weighted representation of the input text using TF-IDF Sparck Jones (1972). To achieve optimal performance, we used cross-validation and GridSearch to rigorously explore the best hyperparameters. Additionally, the best-performing classifiers and hyperparameters varied across tasks, reflecting the distinct nature of each task. Consequently, we conducted task-specific hyperparameter searches for each task respectively.
@@
PLM-based. We first investigated BERT and its various variants for the tasks of answer relevance and answer readability, as these are more challenging compared to the first two tasks. Then, we explored BERT-based models specifically trained on financial text data, such as financial news, research reports, and financial encyclopedia entries (FinBERT, [^4] https://github.com/valuesimplex/FinBERT Mengzi-fin (Zhang et al., 2021), and DKPLM (Zhang et al., 2022)). Furthermore, we examined whether fine-tuned PLMs on specific tasks, such as NLI (Natural Language Inference) and QA (Question Answering), would offer any benefits for these tasks (SBERT-nli (Reimers and Gurevych, 2019) and RoBERTa-extractive-qa (Liu et al., 2019)). Finally, we also examined the performance of a large BERT model (BERT (Large)). [^5] https://huggingface.co/yechen/bert-large-chinese Notably, unless otherwise specified, all PLMs employed in this study are the base and Chinese versions. Links for these PLMs can be found in Supplement Section D.
@@
PLM with CPT-based. Given the prevalence of abbreviations and specialized terms in financial texts, general domain PLMs often struggle with comprehension, especially for the challenging task of answer relevance. To alleviate this issue, we curated a substantial volume of domain-specific texts and conducted continued pre-training on general PLMs. Specifically, we collected over 900,000 pairs of Q&A from the SSE and SZSE platforms by web scraping. High-performing models from previous experiments were then selected for continued pretraining via masked language modeling. This pre-training on a large-scale, domain-specific dataset enhances the PLM’s ability to understand the financial context, capture more specialized information, and improve performance in subsequent downstream tasks.
@@
LLM-based. We further evaluate advanced LLMs in the most challenging task of answer relevance. Different from previous methods, the LLM-based method leverages LLMs’ in-context learning ability to perform the task with zero-shot instead of specific fine-tuning. We select the most advanced LLM (GPT-4) and instruct it with task description and evaluation criteria, requiring the model to directly output relevance score (Direct score). Moreover, we incorporate the Chain of Thought (CoT) (Wei et al., 2022) technique, which instructs the GPT-4 to provide specific reasoning along with the relevance scores (Score with CoT). Prompt and details are listed in Supplement Section G.

### v2

@@ Abstract
Accurate and transparent financial information disclosure is crucial in the fields of accounting and finance, ensuring market efficiency and investor confidence. Among many information disclosure platforms, the Chinese stock exchanges’ investor interactive platform provides a novel and interactive way for listed firms to disclose information of interest to investors through an online question-and-answer (Q&A) format. However, it is common for listed firms to respond to questions with limited or no substantive information, and automatically evaluating the quality of financial information disclosure on large amounts of Q&A pairs is challenging. This paper builds a benchmark FinTruthQA, that can evaluate advanced natural language processing (NLP) techniques for the automatic quality assessment of information disclosure in financial Q&A data. FinTruthQA comprises 6,000 real-world financial Q&A entries and each Q&A was manually annotated based on four key evaluation criteria: question identification, question relevance, answer readability, and answer relevance. We benchmarked various NLP techniques on FinTruthQA, including statistical machine learning models, pre-trained language model and their fine-tuned versions, as well as large language models (LLMs). Experiments showed that existing NLP models have strong predictive ability for question identification and question relevance tasks, but are suboptimal for answer readability and answer relevance tasks. By establishing this benchmark, we provide a robust foundation for the automatic evaluation of information disclosure, significantly enhancing the transparency and quality of financial reporting. FinTruthQA can be used by auditors, regulators, and financial analysts for real-time monitoring and data-driven decision-making, as well as by researchers for advanced studies in accounting and finance, ultimately fostering greater trust and efficiency in the financial markets.
@@ 1 Introduction
Within the financial market, enhancing transparency and timeliness in information disclosure can mitigate information friction, empower investors to make informed decisions, and foster market efficiency. According to Lee & Zhong (2022), retail investors struggle to integrate the existing public information, and increased online platform activities are associated with more active market transactions, such as greater trading volume, return volatility, and market liquidity. Novel investor interactive platforms (such as online forums established by the Shanghai Stock Exchange (SSE) and Shenzhen Stock Exchange (SZSE) ) offer a unique avenue for retail investors to engage directly with listed firms. In contrast to conventional modes of disclosure, such as mandatory disclosures and traditional voluntary disclosures like conference calls, which follow specific procedures and formats, these novel online platforms enable direct, interactive, near real-time communication between firms and ordinary investors through the format of Q&A. By analyzing the interactions on these platforms, researchers can gain insights into how effectively companies communicate with investors, the types of information that investors most value, and how the disclosure practices of listed firms may impact their corresponding trading patterns in financial markets. This will enhance financial market regulation and provide deeper insights into corporate disclosure practices.
@@
The interactive nature of these platforms provides a rich dataset for assessing the quality of information disclosure. However, the sheer volume of data and the diversity in the quality of questions and responses pose significant challenges for large-scale analysis. Leuz & Wysocki (2016) summarizes that the commonly used measures of disclosure in the financial sector primarily capture the quantity rather than the quality of the information provided. However, it is possible to construct the quality dimension of disclosure by focusing on specific disclosures, and Natural Language Processing (NLP) emerges as a promising solution to evaluate the quality of information disclosure at scale. Existing applications of NLP in automatic financial and accounting text analysis have demonstrated success in various domains, such as automatic analysis of annual financial reports(Smailović et al., 2017) and automatic sentiment analysis system on financial news articles(Van de Kauter et al., 2015). However, to the best of our knowledge, there is no standard dataset that evaluates the quality of financial information disclosure on these Q&A platforms.
@@
In this work, we present a Q&A information disclosure quality assessment benchmark that evaluates Chinese Q&A data sourced from interactive platforms for communication between investors and listed companies, as established by SSE and SZSE. In developing this model, we have explored several technical challenges, including the interpretation of natural language, context sensitivity, and the subjective nature of relevance and readability assessments. This paper details the development, implementation, and evaluation of our NLP model, emphasizing its potential impact on the financial sector and the broader implications for automated quality assessment in investor communications. Datasets and code are available at https://anonymous.4open.science/r/FinTruthQA-218F.
@@ 2 Related work
The measurement issue pertaining to disclosures has been a persistent challenge in empirical accounting research(Leuz & Wysocki, 2016). Predominantly, comprehensive measures have focused on quantifying disclosures, encompassing criteria such as the presence or absence of quarterly and annual reports, as well as numerical counts of activities like conference calls and earnings forecasts(Lee & Zhong, 2022). Another commonly used measure is the Association for Investment Management and Research (AIMR) rankings, based on annual surveys of financial analysts who evaluate U.S. firms’ disclosure practices(Nagar et al., 2003; Healy et al., 1999; Welker, 1995). However, they are limited to large U.S. firms and cover only a specific time period. Furthermore, alternative measures designed for specific disclosures typically employ a keyword-based approach to analyze the tone and sentiment of well-established, traditional disclosures such as earnings forecasts and conference call transcripts(Bochkay et al., 2023; Frankel et al., 2010; Li, 2010).
@@
Financial documents, such as annual reports, financial news articles, and economic policy statements, are typically composed in technical language in a non-structured text format and possess unique features distinct from other domains.(Chang et al., 2016) Recent studies have focused on developing pre-trained models tailored for the financial domain (Araci, 2019; Yang et al., 2020; Liu et al., 2021), with a primary emphasis on the English language. These efforts are predominantly oriented towards sentiment classification tasks in the financial context (Wan et al., 2021). Additional research in this field has concentrated on question-answering tasks, involving datasets that combine textual and tabular information, where numerical analysis is often essential to derive answers (Zhu et al., 2021; Zhao et al., 2022; Li et al., 2022). To our knowledge, there has been no prior research or dataset dedicated to the automatic quality assessment of financial Q&A data.
@@ 3.1 Data collection
The data for this study was obtained from the interactive platforms for communication between investors and listed companies established by the Shanghai Stock Exchange (Shangzhen E-Hudong [^1] https://sns.sseinfo.com/) and the Shenzhen Stock Exchange (Hudongyi  [^2] https://irm.cninfo.com.cn/). These two stock exchange sites cover almost all listed companies in China. These Q&A interactions are authentic, generated by human users and company representatives, and are regulated by the China Securities Regulatory Commission, ensuring the accuracy of disclosures. The companies involved are legally obligated to ensure the accuracy of their disclosures, and failure to do so could result in legal consequences. This regulatory oversight ensures that the data we used reflects genuine interactions. We collected over 900,000 Q&A entries, among which we randomly selected 6,000 samples for manual annotation, with 3,000 entries from each exchange’s platform. The data collection involved using web scraping techniques to extract the Q&A entries from the official platforms. The data were then manually reviewed and processed to ensure quality and consistency. The process was overseen by a PhD student from the Business School of XXX University. The ethics statement, including considerations of general ethical conduct and potential negative societal impacts, can be found in Appendix Section A. We also give several examples in Appendix Section G to demonstrate the wide range of investor inquiries covered in our dataset.
Figure 2: Example of a Q&A pair and four quality evaluation tasks.
3.2 Data pre-processing
We found that the Shanghai Stock Exchange (SSE) data includes the company name and stock code at the beginning of the question texts, while the Shenzhen Stock Exchange (SZSE) data does not. To assess the impact of this discrepancy, we conducted experiments comparing Q&A pairs with and without the company name as a prefix and found negligible differences. Therefore, we removed these prefixes from the SSE data in our experiments.
@@ 3.3 Data annotation
We focus on four key information disclosure quality evaluation criteria [^3] In this paper, the terms ”criteria” and ”tasks” are used interchangeably. ”Criteria” refers to the specific standards used to evaluate the quality of the Q&A pairs, while ”tasks” refers to the individual evaluation processes performed based on these criteria, particularly in the experimental sections.: question identification, question relevance, answer relevance, and answer readability. These criteria are widely recognized as crucial indicators of information quality and are important for investors to consider when evaluating Q&A information. Specifically, the criteria in FinTruthQA are defined as follows:
@@
    Question identification: This criterion assesses whether the question from an investor is a genuine question, as opposed to a suggestion or complaint. Each question is assigned a binary label: Positive, indicating it is a real question, or Negative, indicating it is not. Data entries with questions labeled as Negative are excluded from further evaluation under the remaining criteria.
@@
    Question relevance: This criterion evaluates whether a question is pertinent to the investment topic, labeling it as Positive if relevant and Negative if not. It assesses whether the question is related to the company’s shareholders, financial indicators, industry conditions, or other investment-related topics.
@@
    Answer readability: This criterion evaluates how easily an answer can be understood and whether it is well-written. The readability score ranges from 1 to 4, with 4 indicating the highest level of readability—clear, concise, and free of ambiguity. A score of 3 indicates moderate readability with some issues, while a score of 2 reflects significant readability challenges. A score of 1 suggests the answer is unreadable, with major ambiguities, incoherent sentences, or content unrelated to the question.
@@
    Answer relevance: This criterion evaluates whether an answer is relevant and informative in response to the question. It assesses whether the answer provides sufficient detail and explanation to address the question. Answers are classified into three levels: Level 1 indicates the answer is completely unrelated to the question (none of the questions is answered); Level 2 indicates the answer is partially related (some questions are answered or it is difficult to judge) ; and Level 3 indicates the answer fully responds to the question (all questions are answered).
The importance of these criteria lies in their collective ability to enhance the quality and reliability of information on investor communication platforms. The annotation process was carried out in three phases: pre-annotation, formal annotation, and data verification and integration. In the pre-annotation phase, we recruited six students from the Accounting Department at XXX University and conducted five rounds of annotation. After the last round, the inter-annotator agreement (IAA) rate was calculated, and the four students with the highest average IAA rates (85.52%)—two undergraduates and two master’s students—were selected for the formal annotation phase. To ensure annotation quality, the four selected annotators were evenly divided into two groups, with each group annotating the same data instances. Discrepancies were first discussed and resolved within the group, with unresolved issues escalated to the project leader for final decisions. After the formal annotation, the project leader verified the data to ensure alignment with the latest annotation criteria and completeness, before integrating the data and summarising key information, such as sample size and score distribution. Figure 2 illustrates an example of the annotation. A detailed description of the annotation process is listed in the Appendix Section B.
@@ 3.4 Quality control
For the answer relevance task, we observed that the raw annotations contained noise due to the subjective nature of this task. To address this, we introduced an additional quality control step to correct the raw annotations. We implemented an enhanced correction process inspired by Confident Learning (Northcutt et al., 2021a) principles. Sample data points demonstrating inaccuracies, along with implementation details, can be found in Appendix Section C.
@@ 3.5 Dataset statistics
The complete annotated dataset comprises 6,000 samples labeled for question identification, question relevance, answer relevance, and answer readability tasks, with each data point having four distinct labels corresponding to these tasks. The dataset was randomly divided into training, validation, and testing sets in a ratio of 4:1:1. The label distributions for each task are detailed in Tables 1. The provided statistics reflect the data after undergoing the quality control process. Note that our later analysis focuses exclusively on real questions. Only for the first task (question identification), we used the full set of 6,000 samples. For the remaining three tasks, we removed non-question samples, as they could skew the model’s understanding and performance, especially for tasks like question relevance, where the context of a real question is crucial for accurate assessment. It is worth noticing that the question relevance labels are highly imbalanced, with only 6 and 7 negative labels in the validation and test sets, respectively. We retained this task as the imbalance reflects the real-world data distribution, and the concept of question relevance is important in financial information disclosure. Figure  3 shows the length distribution of questions and answers, measured in characters, for all 6,000 QA pairs.
@@
Task Label # Train # Valid # Test # Label Freq. Total
Question identification Negative 327 90 81 498 8.30% 6,000
Positive 3,673 910 919 5,502 91.70%
Question relevance Negative 714 6 7 27 0.49% 5,502
Positive 3,654 911 910 5,475 99.51%
Answer readability 1 106 190 198 1,094 19.88% 5,502
@@ 4.1 Model selection
To fully evaluate the performance of different methods in FinTruthQA, we benchmarked several models that cover Machine Learning (ML)-based, Pre-trained Language Model (PLM)-based, PLM with continued Pre-Training (CPT)-based, and LLM-based approaches. Given the distinct nature and varying difficulty of the tasks, we applied tailored approaches to each accordingly.
@@
ML-based. We modeled these tasks as text classification problems and employed various advanced machine learning models, including Logistic Regression (LR)(Hosmer Jr et al., 2013), Support Vector Machine (SVM)(Hearst et al., 1998), K-nearest neighbor (KNN)(Peterson, 2009), and Random Forest (RF)(Breiman, 2001). For semantic feature extraction, we first utilized the Jieba toolkit  [^4] https://github.com/fxsjy/jieba for Chinese word segmentation and then extracted weighted text representations using TF-IDF Sparck Jones (1972). To achieve optimal performance, we used cross-validation and GridSearch to rigorously explore the best hyperparameters. The best-performing classifiers and hyperparameters varied across tasks, reflecting the distinct nature of each task. As a result, we performed task-specific hyperparameter searches for each task individually.
@@
PLM-based. In our experiments, we found that the tasks of question identification and question relevance were relatively straightforward, with BERT achieving excellent performance on these tasks. Therefore, our focus shifted to exploring the effectiveness of PLM-based methods on the more challenging tasks of answer relevance and answer readability. Specifically, we investigated a range of BERT-based models, including those specifically trained on financial text data, such as financial news, research reports, and financial encyclopedia entries (FinBERT, [^5] https://github.com/valuesimplex/FinBERT Mengzi-fin (Zhang et al., 2021), and DKPLM (Zhang et al., 2022)). Furthermore, we examined whether fine-tuned PLMs on specific tasks, such as NLI (Natural Language Inference) and QA (Question Answering), would offer any benefits for these tasks (SBERT-nli (Reimers & Gurevych, 2019) and RoBERTa-extractive-qa (Liu et al., 2019)). Finally, we also examined the performance of a large BERT model (BERT (Large)). [^6] https://huggingface.co/yechen/bert-large-chinese Notably, unless otherwise specified, all PLMs employed in this study are the base and Chinese versions. Links for these PLMs can be found in Appendix Section D.
@@
PLM with CPT-based. General domain PLMs, trained on broad texts, often lack the specialized knowledge required to handle the nuanced language and terminology in finance. This limitation is especially pronounced in tasks like answer relevance, where understanding domain-specific jargon and abbreviations is essential for accurate predictions. To bridge this gap, we curated a substantial volume of domain-specific texts and conducted continued pre-training on general PLMs. Specifically, we collected over 900,000 Q&A pairs from the SSE and SZSE platforms by web scraping. The top-performing models from previous experiments were then selected for continued pre-training using masked language modeling (MLM). This pre-training on a large-scale, domain-specific dataset enhances the PLM’s ability to understand financial contexts, capture more specialized information, and achieve better performance in downstream tasks.
@@
LLM-based. We further evaluate some advanced LLMs, including GPT-4 (Achiam et al., 2023), LLaMA-3.1 (Dubey et al., 2024), Qwen2 (Yang et al., 2024), and Mistral (Jiang et al., 2023), on the most challenging answer relevance task. Due to their in-context learning ability, LLMs can perform this task in a zero-shot manner, without requiring specific fine-tuning. Specifically, we provided LLMs with task descriptions and evaluation criteria, instructing them to directly output relevance scores (Direct score). Moreover, we incorporated the Chain of Thought (CoT) (Wei et al., 2022) technique, prompting the LLMs to provide specific reasoning along with the relevance scores (Score with CoT). Full prompts and details are provided in Appendix Section F.